Let X and Y be two independent random variables having joint probability density function f(x, y) = 1/ 2πσ2 e − (x−µ) 2 σ2 e − (y−µ) 2 σ2 − ∞ < x, y < ∞

Find the moment generating function of Z = X+Y 2 and hence the mean and variance of Z


1
Expert's answer
2022-04-04T16:20:40-0400

"f\\left( x,y \\right) =\\frac{1}{2\\pi \\sigma ^2}\\exp \\left( -\\frac{\\left( x-\\mu \\right) ^2}{2\\sigma ^2} \\right) \\exp \\left( -\\frac{\\left( y-\\mu \\right) ^2}{2\\sigma ^2} \\right) \\\\This\\,\\,is\\,\\,a\\,\\,density\\,\\,of\\,\\,i.i.d. random\\,\\,variables\\,\\,X,Y\\sim N\\left( \\mu ,\\sigma ^2 \\right) \\\\M_Z\\left( t \\right) =Ee^{tZ}=E^{t\\left( X+Y^2 \\right)}=\\\\=Ee^{tX}Ee^{tY^2}=\\exp \\left( \\mu t+\\frac{1}{2}\\sigma ^2t^2 \\right) \\int_{-\\infty}^{+\\infty}{\\frac{1}{\\sqrt{2\\pi \\sigma ^2}}\\exp \\left( ty^2-\\frac{\\left( y-\\mu \\right) ^2}{2\\sigma ^2} \\right)}dy=\\\\=\\exp \\left( \\mu t+\\frac{1}{2}\\sigma ^2t^2 \\right) \\int_{-\\infty}^{+\\infty}{\\frac{1}{\\sqrt{2\\pi \\sigma ^2}}\\exp \\left( -\\frac{\\left( y-\\frac{\\mu}{1-2t\\sigma ^2} \\right) ^2}{\\frac{2\\sigma ^2}{1-2t\\sigma ^2}} \\right) \\exp \\left( \\mu ^2t \\right) dt}=\\\\=\\left[ \\begin{array}{c}\tf\\left( t \\right) =\\frac{1}{\\sqrt{2\\pi \\frac{\\sigma ^2}{1-2t\\sigma ^2}}}\\exp \\left( -\\frac{\\left( y-\\frac{\\mu}{1-2t\\sigma ^2} \\right)}{\\frac{2\\sigma ^2}{1-2t\\sigma ^2}} \\right) \\sim N\\left( \\frac{\\mu}{1-2t\\sigma ^2},\\frac{\\sigma ^2}{1-2t\\sigma ^2} \\right) \\Rightarrow\\\\\t\\Rightarrow \\int_{-\\infty}^{+\\infty}{f\\left( t \\right) dt}=1\\\\\\end{array} \\right] =\\\\=\\exp \\left( \\mu t+\\frac{1}{2}\\sigma ^2t^2 \\right) \\frac{\\exp \\left( \\mu ^2t \\right)}{\\sqrt{1-2t\\sigma ^2}}=\\frac{\\exp \\left( \\left( \\mu +\\mu ^2 \\right) t+\\frac{1}{2}\\sigma ^2t^2 \\right)}{\\sqrt{1-2t\\sigma ^2}}\\\\EZ=\\frac{d}{dt}M_Z\\left( t \\right) |_{t=0}=\\\\=\\exp \\left( \\left( \\mu +\\mu ^2 \\right) t+\\frac{1}{2}\\sigma ^2t^2 \\right) \\frac{\\left( -2\\left( \\mu +\\mu ^2 \\right) \\sigma ^2t+\\mu +\\mu ^2-2\\sigma ^4t^2+\\sigma ^2t+\\sigma ^2 \\right)}{\\left( 1-2t\\sigma ^2 \\right) ^{3\/2}}|_{t=0}=\\\\=\\mu +\\mu ^2+\\sigma ^2\\\\EZ^2=\\frac{d^2}{dt^2}M_Z\\left( t \\right) |_{t=0}=\\\\=\\exp \\left( \\left( \\mu +\\mu ^2 \\right) t+\\frac{1}{2}\\sigma ^2t^2 \\right) \\left( \\frac{3\\sigma ^4}{\\left( 1-2t\\sigma ^2 \\right) ^{5\/2}}+\\frac{2\\sigma ^2\\left( \\mu +\\mu ^2+\\sigma ^2t \\right)}{\\left( 1-2t\\sigma ^2 \\right) ^{3\/2}}+\\frac{\\left( \\mu +\\mu ^2+\\sigma ^2t \\right) ^2+\\sigma ^2}{\\left( 1-2t\\sigma ^2 \\right) ^{1\/2}} \\right) |_{t=0}=\\\\=3\\sigma ^4+2\\sigma ^2\\left( \\mu +\\mu ^2 \\right) +\\left( \\mu +\\mu ^2 \\right) ^2+\\sigma ^2\\\\Var\\left( Z \\right) =3\\sigma ^4+2\\sigma ^2\\left( \\mu +\\mu ^2 \\right) +\\left( \\mu +\\mu ^2 \\right) ^2+\\sigma ^2-\\left( \\mu +\\mu ^2+\\sigma ^2 \\right) ^2=\\\\=\\sigma ^2+2\\sigma ^4"